OpenStudy (anonymous):

find two positive numbers whose product is 800 and the sum of the first plus twice the second is a minimum?

OpenStudy (amistre64):

sounds like a lagrange multiplier problem

OpenStudy (amistre64):

ab=800 Min=a+2b hmm

OpenStudy (amistre64):

f(a,b,L) = a+2b+L(ab-800) maybe, then partial it out

OpenStudy (anonymous):

still dont get it :(

OpenStudy (amistre64):

fa = 1+Lb = 0 b = -1/L fb = 2+La = 0 a = -2/L fL = ab-800 = 0 ; sub in a and b to solve for L (-1/L * -2/L) - 800 = 0 2/L^2 - 800 = 0 (2 - 800L^2)/L^2 = 0 2 - 800L^2 = 0 ; when L^2=2/800

OpenStudy (anonymous):

its about derivatives of exponential fuction.

OpenStudy (amistre64):

it is? well, lets see if this works out cause I need to practice my Lagrange multipliers anyhoos :)

OpenStudy (anonymous):

what is L?

OpenStudy (amistre64):

L^2=2/800 L = -sqrt(2/800) b = 1/sqrt(2/800) = 20 a = 2/sqrt(2/800) = 40 maybe

OpenStudy (anonymous):

i mean L = minimum?

OpenStudy (amistre64):

L is a Lagrange multiplier; which if you havent learnt it yet, you will later on

OpenStudy (amistre64):

Its a way to compute a system of equations that have constraints

OpenStudy (anonymous):

o... i haven't learned that one yet.. hahah is there any other way solving this problem by either using Derivatives of Exponential Fuction or Derivatives of Log?

OpenStudy (amistre64):

i dont think there is a way to do it with exp or logs; but its been awhile so I could just be forgetting stuff

OpenStudy (amistre64):

id solve it as a system of equations and derive them to see where the minimum is

OpenStudy (amistre64):

ab=800; solve for say a=800/b f(a,b)=a+2b; plug in our a value 800/b + 2b ; and derive -800/b^2 + 2 i think is the derivative

OpenStudy (amistre64):

when that equals 0 we got critical points

OpenStudy (amistre64):

-800+2b^2 = 0 when b^2 = 800/2 b = sqrt(400) = 20

OpenStudy (amistre64):

ab=800 a(20) = 800 a = 800/20 a = 40, jut like the lagrange worked out too yay!!

OpenStudy (anonymous):

800/b + 2b ; and derive -800/b^2 + 2 i think is the derivative whats derive and how did you do this?

OpenStudy (amistre64):

derive means to "take the derivative"

OpenStudy (amistre64):

i assumed you knew that since you asked about taking the derivatives of exponential and log functions

OpenStudy (anonymous):

we just learned that haha thank you !

OpenStudy (anonymous):

but what is the meaning of the 'minimum?'

OpenStudy (amistre64):

minimum, least, lowest point, can get lower than this, bottom of the barrel, etc ....

OpenStudy (amistre64):

the smallest number

OpenStudy (anonymous):

so in this question, it means the smaller critical point yes?

OpenStudy (amistre64):

a+2b 40 + 2(20) = 80 80 is the smallest number you can get with the factors of 800

OpenStudy (amistre64):

yes

OpenStudy (amistre64):

the critical points themselves need to be dbl chked to make sure they are mins or maxs

OpenStudy (anonymous):

got it! thank youuuu ;)